Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎


Crew 100A Log Book for February 17, 2011

Photos of the day

Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA #5: 15:30 - 17:40

Brief Narrative of Field Mission Results:

We continued with the habitability experiments during the morning.

Sound and smell experiments were the tests that Ayako set for some of the crew members. The day was mainly spent installing the equipment necessary to extend the range of the wireless network to be able to control the rover and

the drone for longer distances. There was an EVA on a real search and rescue mission to the parrot UAV that flew away yesterday to the north of the habitat. We tried hard to find it for more than 2 hours but our effort was not rewarded yet. There were a lot of lessons learned that will serve to enhance future EVA missions of this type.

The rest of the day was spent preparing tomorrow's EVA that will include a backup Parrot UAV, a Rover and 3 Humans for a science mission.

EVA Data/Interpretations: On a search mission it is important to pinpoint the different areas to be searched by the different crew members. If this does not happen there might be overlapping and waste of resources.

Engineering/Hab Maintenance: N/A

Plans for Tomorrow: Continue with experiments: habitability and stress communication EVA. EVA to study symbiotic interaction between UAV, rover and humans.

Reports scheduled for today: Engineering, Commander, Journalist and EVA#5

Inventory: None

Support Requested: N/A

Miscellaneous: We received the packages for the experiments of crew 100B.

Commander's Report

Ivo Ferreira Reporting

Gone with the wind.

Ayako greeted us with the breakfast followed by some other experiments to study habitability by exposing us to nice smells, colors or even some relaxing sounds like the water flowing down a river, reminded us of a shower...  a shower that we are already starting to miss...

While some of us used the time to work a bit on our different experiments, Jeffrey was contemplating his "newer babies" that will help extend the range to teleoperate the UAV or the Rover to distances up to 7 km. Maybe we can send the Rover to pick a Pizza for us in Hanksville!

After lunch, me, Romain and Audrey set out on an EVA to search and Rescue the UAV that flew away. By looking at the last transmission of the rover and knowing its range we were able to determine the search area, and planned the best strategy to cover it. Unfortunately, this area was in the order of a few square miles so it was like finding a needle in an Haystack. After 2 hours, we had our share of rock

climbing and eating dust and decided to get back to the Habitat.

The rest of the day was spent preparing a complex EVA that we will conduct tomorrow using the UAV, the rover, and Humans, simultaneously.

Let's hope that tomorrow the EVA shows a symbiotic relationship between the different actors and that the other Parrot does not decide to leave as well.

Engineering Report

Audrey Lan San Reporting


· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good

Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.56

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.58

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 8 Gallons



· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):

Water Status:

· Outside Potable Water Tank Level (inches from bottom): 9

· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full

· Hab Potable Water Tank Level (inches from bottom): 9

· Potable Water Meter Reading: 422107



· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments:

Photographs will be sent with the next mail.


· Telescope Used? (Yes/no) No

Observatory Notes:



· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123611.6

· Notes/Comments on Hab Car:

ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: Yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:

ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed

ATV 3 (Viking I):

· Used: Yes

· Oil Checked: Yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:

ATV 4 (Viking II):

· Used: yes

· Oil Checked: yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:

Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 65

· Actual temperature upstairs (Farenheit): 72

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 64

Computers/Networking Infrastructure:

Notes/comments on internet/computers?: All is well

General Engineering

1. Notes/Comments:

Nothing special happened

EVA 5 Report

Romain Benchenafi Reporting

EVA #5, 15:30-17:40

Crew members : Ivo FERREIRA, LAN SAN Audrey,  BENCHENAFI Romain

Site location : UTM 518140, 4251346

Transit mode : ATVs

Objectives : Search and rescue the UAV thanks to the last data received from it.

Results : Managed to coordinate the team to search efficiently in the designated area. UAV still missing.

Lesson learned :  weather conditions can make radio conversations hard to understand.

Journalist Report

Romain Benchanafi Reporting

Today we still had to face hard windy conditions. We have experimented to find the maximum range of the wifi network of our second UAV in order to estimate the area in which the lost drone is likely to be.

The crew has been working hard this day searching for the UAV over  a quite wide space. This was an exhausting and unsuccessful mission, consequently some entertainment would be needed tonight.

Cook's Report

Kévin Theatre Reporting

1a) Today was:

A cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?


2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Mixed pepper,onions,tomato dices,carrots, chili, ham, couscous.

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)  


4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.  


5)  Lessons learned  (if any)

Ham is horrible

6) Comments/questions for Kim and Jean, the food study investigators